A consumer advocacy group suspects that a local supermarket’s 500 grams of sugar actually weigh less than 50 grams. The group look a random sample of 20 such packages, weigh each one, and found the mean weight for the sample to be 496 grams with standard deviation of 8 grams. Using 1% significance level, would you conclude that the mean weight is less than 500 grams? Also, find the 99% confidence interval of the true mean.


1
Expert's answer
2022-05-09T17:56:26-0400

We need to construct the "99\%" confidence interval for the population mean "\mu."

The critical value for "\alpha = 0.01" and "df = n-1 = 19" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.860935."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}},\bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(496-2.860935\times\dfrac{8}{\sqrt{20}},"

"496+2.860935\times\dfrac{8}{\sqrt{20}})"

"=(490.8822,501.1178)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "490.8822 < \mu < 501.1178," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(490.8822,501.1178)."


The following null and alternative hypotheses need to be tested:

"H_0:\mu\ge500"

"H_a:\mu<500"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=19" degrees of freedom, and the critical value for a left-tailed test is "t_c = -2.539483."

The rejection region for this left-tailed test is "R = \{t: t < -2.539483\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{496-500}{8/\sqrt{20}}\approx-2.236068"

Since it is observed that "t = -2.236068 \ge -2.539483=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value for left-tailed test, "df=19" degrees of freedom, "t=-2.236068" is "p = 0.01877," and since "p = 0.01877 \ge 0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 500, at the "\alpha = 0.01" significance level.